OpenStudy (anonymous):

City A and City B had two different temperatures on a particular day. On that day, four times the temperature of City A was 8° C more than 3 times the temperature of City B. The temperature of City A minus twice the temperature of City B was −3° C. What was the temperature of City A and City B on that day?

OpenStudy (anonymous):

City A was 5° C, and City B was 4° C. City A was 3° C, and City B was −1° C. City A was 8° C, and City B was −3° C. City A was 5° C, and City B was −5° C

OpenStudy (anonymous):

would it be the second answer

OpenStudy (welshfella):

let temperature in CitY A be x and temperature in City B be y

OpenStudy (welshfella):

4x = 3x + 8 x - 2y = -3

OpenStudy (anonymous):

wait so why would it be 4x instead of 8x

OpenStudy (welshfella):

you can find which option is correct by plugging in the given values and see which ones fit

OpenStudy (welshfella):

4 times x = 4x

OpenStudy (welshfella):

four times the temperature of city A = 4 times x = 4x

OpenStudy (anonymous):

ok thank you that makes alot more sense

OpenStudy (welshfella):

you can solve the system of equations or as I said earlier you can plug in the values in the options to see which fit.

OpenStudy (anonymous):

ill plug in the values

OpenStudy (anonymous):

4(3)=3(3)+8

OpenStudy (welshfella):

first try x=5 and y = 4 4(5) = 3(4) + 8 20 = 12 + 8 - those fit

OpenStudy (welshfella):

now plug same values into x - 2y = -3

OpenStudy (anonymous):

ok gotcha

OpenStudy (anonymous):

(5)-2(4)=-3

OpenStudy (anonymous):

5-8=-3 -3=-3

OpenStudy (anonymous):

it works

OpenStudy (welshfella):

right

OpenStudy (anonymous):

thanks